# BoundMatch: Boundary detection applied to semi-supervised segmentation for urban-driving scenes

### v3

@@ E.1 BDD100K
We show the qualitative results on the BDD100K dataset in Fig. 13. Compared to UniMatch, our SAMTH+BoundMatch generally outputs more accurate segmentations. For example, in the first and last rows, the bus and truck is accurately segmented with BoundMatch. In the second and fifth rows, the road and buildings are accurately segmented with BoundMatch, where UniMatch fails to do so. The fourth row also shows BoundMatch’s robustness against sun flares and shadows, being able to segment the road more accurately. However, there remains some challenges when the dataset’s labels are ambiguous; such as the “sidewalk” region next to the road might be considered a “vegetation” in the first row.
E.2 Pascal VOC 2012
@@
We show the qualitative results on the Pascal VOC 2012 dataset in Fig. 14. This dataset is especially challenging due to the presence of small objects and complex backgrounds, but our SAMTH+BoundMatch method demonstrates improved segmentation quality over SAMTH especially around object boundaries. For example, the first row shows that BoundMatch is able to delineate the airplane and have better boundary quality compared to without using boundary consistency regularization. Also, the third row also suggests that the model learns to separate the category of interest from the background more accurately.
E.3 DINOv2-S
@@
We show the qualitative results on the Cityscapes dataset using DINOv2-S pretrained ViT in Fig. 15. DINOv2-S is a strong foundation model that has been pre-trained on a large dataset, and it is known to perform well on various vision tasks. As shown in the samples, BoundMatch is able to improve the segmentation quality compared to SAMTH. For example, the first row shows that BoundMatch is able to delineate the road/sidewalk and the car around their boundaries more accurately. Far away objects such as the wall and the thin poles are also better segmented with BoundMatch. Other samples also provide similar observations.

### v4

@@ E.2 BDD100K
    We show the qualitative results on the BDD100K dataset in Fig.˜15. Compared to UniMatch, our SAMTH+BoundMatch generally outputs more accurate segmentations. For example, in the first and last rows, the bus and truck is accurately segmented with BoundMatch. In the second and fifth rows, the road and buildings are accurately segmented with BoundMatch, where UniMatch fails to do so. The fourth row also shows BoundMatch’s robustness against sun flares and shadows, being able to segment the road more accurately. However, there remains some challenges when the dataset’s labels are ambiguous; such as the “sidewalk” region next to the road might be considered a “vegetation” in the first row.
    E.3 Pascal VOC 2012
@@
    We show the qualitative results on the Pascal VOC 2012 dataset in Fig.˜16. This dataset is especially challenging due to the presence of small objects and complex backgrounds, but our SAMTH+BoundMatch method demonstrates improved segmentation quality over SAMTH especially around object boundaries. For example, the first row shows that BoundMatch is able to delineate the airplane and have better boundary quality compared to without using boundary consistency regularization. Also, the third row also suggests that the model learns to separate the category of interest from the background more accurately.
    E.4 DINOv2-S
@@
    We show the qualitative results on the Cityscapes dataset using DINOv2-S pretrained ViT in Fig.˜17. DINOv2-S is a strong foundation model that has been pre-trained on a large dataset, and it is known to perform well on various vision tasks. As shown in the samples, BoundMatch is able to improve the segmentation quality compared to SAMTH. For example, the first row shows that BoundMatch is able to delineate the road/sidewalk and the car around their boundaries more accurately. Far away objects such as the wall and the thin poles are also better segmented with BoundMatch. Other samples also provide similar observations.
    Appendix F BoundMatch in other domains
    In Tab.˜19, we show the results on the ACDC dataset [68] which is a medical image segmentation dataset. We follow the same experimental setup as in [9] and use UNet as the base model. To integrate BoundMatch, we took the hierarchical features from the encoder to the boundary detection head follwing Sec.˜B.1. We found that BoundMatch does not improve the performance of UniMatch on this dataset. This could be due to the fact that our architectures and losses for the introduced BoundMatch would need to be further optimized for medical images, which is out of the scope of this paper.
    In Tab.˜20, we show the results on the LoveDA dataset [69] which is a remote sensing image segmentation dataset. We use DeepLabV3+ as the base model and produced results on $\nicefrac{{1}}{{32}}$ and $\nicefrac{{1}}{{16}}$ splits. We can see that BoundMatch consistently improves the performance of SAMTH and UniMatch on this dataset, providing potentials that BoundMatch can be generalized to other domains.